OpenStudy (anonymous):

How do I solve -x+6+4x=1+5x+3

OpenStudy (anonymous):

the answer is X = 1

OpenStudy (anonymous):

you have to combine like terms then solve.

OpenStudy (alesia/bailey):

-x+6+4x=1+5x+3 you combine like terms, anything you do on one side, you do to the other -x+6+4x=1+5x+3 -x+4x=3x 3x+6=1+5x+3 -3x -3x 6=1+2x+3 6=4+2x -4 -4 2=2x (divide the 2 for 2x, this is because you are trying to isolate x) /2 /2 1=x